# ConceptMaster: Multi-Concept Video Customization on Diffusion Transformer Models Without Test-Time Tuning

### v1

@@ Abstract
Text-to-video generation has made remarkable advancements through diffusion models. However, Multi-Concept Video Customization (MCVC) remains a significant challenge. We identify two key challenges in this task: 1) the identity decoupling problem, where directly adopting existing customization methods inevitably mix attributes when handling multiple concepts simultaneously, and 2) the scarcity of high-quality video-entity pairs, which is crucial for training such a model that represents and decouples various concepts well. To address these challenges, we introduce ConceptMaster, an innovative framework that effectively tackles the critical issues of identity decoupling while maintaining concept fidelity in customized videos. Specifically, we introduce a novel strategy of learning decoupled multi-concept embeddings that are injected into the diffusion models in a standalone manner, which effectively guarantees the quality of customized videos with multiple identities, even for highly similar visual concepts. To further overcome the scarcity of high-quality MCVC data, we carefully establish a data construction pipeline, which enables systematic collection of precise multi-concept video-entity data across diverse concepts. A comprehensive benchmark is designed to validate the effectiveness of our model from three critical dimensions: concept fidelity, identity decoupling ability, and video generation quality across six different concept composition scenarios. Extensive experiments demonstrate that our ConceptMaster significantly outperforms previous approaches for this task, paving the way for generating personalized and semantically accurate videos across multiple concepts.
Figure 1: We propose ConceptMaster, a Multi-Concept Video Customization (MCVC) method that could create high-quality and concept-consistent customized videos based on given multiple reference images without additional test-time tuning. Our ConceptMaster is capable of handling diverse customized scenarios, including but not limited to 1) multiple persons, 2) persons with livings, 3) persons with stuffs, 4) multiple livings, 5) livings with stuffs and 6) persons with both livings and stuffs.
footnotetext: Corresponding author  Work done during an internship at KwaiVGI, Kuaishou Technology
@@ 1 Introduction
Diffusion-based text-to-video generation models, trained on extensive text-video data pairs, have demonstrated remarkable success in generating high-quality videos from textual inputs [15, 6, 58, 3, 71, 64, 56, 52, 41, 26, 18, 11, 2]. These advancements have sparked increasing interest in personalizing video generation through user-defined concepts. Some methods have been proposed to produce customized videos using additional image guidance, with proven effectiveness for the customization on object [25], human [17], style [37], etc.
Existing approaches to concept customization primarily fall into two methodological categories: tuning-based solutions and pretrain-based methods. The tuning-based solutions [50, 13, 30, 16, 9] typically first optimize model parameters (e.g., variants of LoRAs [22] or fully training latent diffusion models [50]) each time for customized concepts and then incorporate them for inference. However, these methods are computationally time-consuming and often require the manual collection of multiple reference samples, rendering them impractical in most time-sensitive and user-friendly scenarios. Conversely, pretrain-based methods [62, 14, 33, 8, 67, 35, 60, 61, 25, 17] aim to integrate visual embeddings into diffusion models at training time in a data-driven manner, enabling personalization without additional test-time parameter tuning. Despite their progress, how to adopt these approaches to simultaneously process multiple concepts in videos to keep both concept fidelity and factorization in a feed-forward approach remains challenging.
@@
To overcome these challenges, we propose ConceptMaster (see Fig. 1), an effective MCVC method that could effectively maintain the fidelity of multiple concepts well and address the identity decoupling problem, even for highly similar concepts. Unlike previous approaches to integrate visual embeddings with textual counterparts [33, 35, 25], or directly aggregate visual embeddings as a whole into diffusion models [67, 61, 17], our key insight is to learn the decoupled multi-concept embeddings and inject into diffusion transformer models in a standalone manner. Specifically, the process includes: 1) Extracting comprehensive visual embeddings from given reference images, where we initially extract dense visual tokens by the CLIP image encoder [45], and integrate a learnable query transformer (Q-Former) network [34] to better represent comprehensive visual embeddings and align with the diffusion model space. 2) Incorporating visual representation with corresponding text description of every concept, where we propose the Decouple Attention Module (DAM) to conduct intra-pair attention to separately bind the extracted visual embedding with corresponding textual embedding for each concept, the process could effectively capture semantic differences across multiple concepts while maintain concept-specific uniqueness. 3) Introducing a novel multi-concept embeddings injection strategy, where we firstly composite the multi-concept embeddings and then inject them using an individual Multi-Concept Injector (MC-Injector), which is a standalone cross-attention layer, into the diffusion transformer models without affecting the original textual cross-attention. This strategy separates the functionality of the original textual cross-attention from the learning process of the newly injected composite multi-concept embeddings, effectively enhancing the representation of multiple identities. The designed ConceptMaster could efficiently create high-fidelity customized videos during inference without additional parameter tuning, which significantly provides the potential for the practicality of real-world applications.
Furthermore, to address the scarcity of suitable MCVC data, we carefully establish a data collection pipeline, which could collect high-quality MCVC data that precisely extract entity images and corresponding text descriptions of diverse concepts in videos. By utilizing this pipeline, we collect over 1.3 million video-entity pairs spanning diverse conceptual domains, including humans, livings, and various object categories. To further facilitate the evaluation, we introduce a Multi-Concept Benchmark (MC-Bench) to comprehensively validate this task from 1) concept fidelity, 2) effectiveness of identity decoupling, and 3) video generation quality across six distinct multi-concept composition scenes. Overall, our key contributions can be summarized:
@@
  • We present a novel strategy of learning the decoupled multi-concept embeddings and injecting into the diffusion models without influencing the original text cross-attention, which guarantees both representation and decoupling of multiple concepts in customized videos.
  • We introduce a dedicated data collection pipeline that enables the creation of high-quality multi-concept video-entity pairs across diverse concepts, and we collect more than 1.3 million suitable MCVC data for the task.
  • We introduce MC-bench, a comprehensive benchmark including six distinct multi-concept composition scenarios for evaluation. We rigorously assess customized videos on concept fidelity, effectiveness of identity decoupling, and video quality. Extensive quantitative and qualitative experiments demonstrate that ConceptMaster significantly outperforms the previous state-of-the-art methods.
@@ 4 ConceptMaster
Figure 3: Overview of the framework of our proposed ConceptMaster.
@@ 4.1 Multi-Concept Video Customization
Problem: Given a caption $T$ describing a video, along with a set of concept images $\{X_{i}|i=1\ldots N\}$ and their corresponding labels $\{Y_{i}|i=1\ldots N\}$ (e.g., a man and a woman with their respective images), where $N$ represents the number of distinct concepts, the task of Multi-Concept Video Customization (MCVC) aims to generate high-quality videos that incorporate all image-defined visual concepts while aligning them with the given descriptive caption $T$. Each concept should maintain its identity as the provided images while precisely expressing its semantic behavior as described in the caption.
@@
4.2 Injecting the Decoupled Multi-Concept Embeddings
@@
Decoupling Intra-Pair Embeddings. After obtaining the appropriate visual representation, we integrate the corresponding text labels to create multi-modal concept representation. While previous works [65, 35] directly combine the visual representation with the corresponding word from the caption embedding $c_{text}=\mathcal{E}_{\text{text}}(T)$, we hope to fully leverage the textual label information associated with related image to enhance the representation specific to each concept. Therefore, unlike these approaches, we employ T5-encoder $\mathcal{E}_{T5}$ to encode each concept label individually to obtain the text representation $\{y_{i}|y_{i}=\mathcal{E}_{T5}(Y_{i}),i=1\ldots N\}$. Subsequently, we introduce the Decouple Attention Module (DAM) to fuse each pair of the visual and text label embedding $\{(x_{i},y_{i})|i=1\ldots N\}$. The DAM operation can be formulated as:
{$&Q_{i}=W_{Q}\cdot x_{i};K_{i}=W_{K}\cdot y_{i};V_{i}=W_{V}\cdot y_{i},\\ &Attention(Q_{i},K_{i},V_{i})=\text{Softmax}(\frac{Q_{i}K_{i}^{T}}{\sqrt{d}})\cdot V_{i},\\ &c_{i}=\text{FFN}(Attention(Q_{i},K_{i},V_{i}))\end{aligned}\right.$ (2)
where $W_{Q}$, $W_{K}$, and $W_{V}$ are projection matrices, $d$ is the embedding dimension, and FFN is a two-layer multi-layer perceptron (MLP) with GLUE [57] as the middle activation function. With the designed DAM, every visual representation could integrate its corresponding textual label to serve as the multi-modal representation for the diffusion transformer models.
@@ 5 Experiments
Methods Concept Fidelity Decoupling Ability Video Quality
@@
CustomDiffusion 27.986 0.592 19.371 0.643 0.531 0.964 13.648 0.557 0.686
SSR-Encoder 24.572 0.598 19.857 0.659 0.517 0.962 13.829 0.518 0.631
IP-Adapter 27.943 0.551 14.829 0.539 0.394 0.977 14.072 0.564 0.701
MS-Diffusion 28.519 0.584 21.060 0.666 0.528 0.976 14.703 0.525 0.704
DreamBooth 18.887 0.517 12.781 0.562 0.44 0.991 8.23 0.499 0.632
ConceptMaster (Ours) 29.461 0.618 22.378 0.678 0.623 0.987 17.585 0.573 0.709
Table 1: Quantitative comparison with different methods on MC-Bench, where bold represents the best result and underline represents the second best result.
@@ 5.1 Experimental Setup
Evaluation Metrics. In order to comprehensively evaluate the methods for the MCVC task, we consider from three different dimensions: 1) Concept fidelity, where we exploit the commonly used CLIP-I and CLIP-T [45] to globally evaluate that if the generated videos match the given reference images and video captions. 2) Decoupling ability, where we firstly utilize OWLv2 [43] to detect the box area of each concept in generated videos, and then we compute CLIP-I and DINO-I [4] scores between the original concept images and specific box areas in generated videos, and CLIP-T between specific box areas and detected labels for local evaluation. 3) Video generation quality, where we adopt the motion smoothness, dynamic degree, aesthetic quality and imaging quality collected by VBench [24] to evaluate the quality of generated videos.
Methods Concept Fidelity Decoupling Ability Video Quality
@@
Merge Textual and Visual Embeddings 29.418 0.569 17.063 0.606 0.367 0.995 3.748 0.549 0.666
IP-Adapter-like 26.812 0.558 17.477 0.609 0.402 0.986 13.946 0.528 0.653
ConceptMaster (Ours) 29.461 0.618 22.378 0.678 0.623 0.987 17.585 0.573 0.709
Table 2: Quantitative comparison of different multi-concept embeddings injection manner on MC-Bench, bold represents the best result and underline represents the second best result. All the methods we compared have been trained on the same data as that used by ConceptMaster.
Methods Concept Fidelity Decoupling Ability Video Quality
@@
Without Q-Former 27.595 0.512 19.256 0.589 0.277 0.98 16.881 0.548 0.631
Without DAM 26.463 0.537 19.718 0.583 0.394 0.985 16.481 0.526 0.665
Concate-MLP 27.576 0.583 20.535 0.643 0.546 0.984 15.125 0.554 0.696
Self-Attn 27.679 0.603 21.946 0.641 0.581 0.986 16.436 0.575 0.705
DAM (Ours) 29.461 0.618 22.378 0.678 0.623 0.987 17.585 0.573 0.709
Table 3: Quantitative comparison of the design choice of Q-Former and DAM modules on MC-Bench, bold represents the best result and underline represents the second best result. All the methods we compared have been trained on the same data as that used by ConceptMaster.
5.2 MC-Bench Evaluation
In order to comprehensively evaluate the performances of the MCVC models on diverse scenarios, we establish a Multi-Concept Benchmark (MC-Bench), where we collect videos from the Internet sources by the same MCVC data collection pipeline mentioned in section 4.3, and we only leave the videos containing multiple concepts, including 1) multiple persons, 2) persons with livings, 3) persons with stuffs, 4) multiple livings, 5) livings with stuffs and 6) persons with both livings and stuffs. The number of samples for each scene is 50, 50, 50, 50, 50, and 33 respectively (Total 283). It is worth being noticed that we carefully guarantee there is no any overlap between the training videos and the constructed MC-Bench.
@@ 5.3 Comparing with other methods
Comparing with naive solutions. We compare several open-sourced multi-concept image customization methods [30, 69, 67, 61], combining with the image-to-video (I2V) generation model I2VGen-XL [68], as a naive solution for the MCVC task with our ConceptMaster. According to the quantitative and qualitative results in Tab. 1 and Fig. 5, we can see that our ConceptMaster has clear advantages on customizing multiple concepts in videos. The naive solution will be subject to the decoupling and representation ability of both two models, the instruction-following capability of I2V models could further influence the quality of generated videos. In contrast, the end-to-end video customization models could achieve better results.
Figure 6: Comparison with tuning-based method DreamBooth.
Comparing with tuning-based methods. We compare the classical tuning-based method DreamBooth [50], where we segment the original videos from MC-Bench into multiple clips for few-shot training examples and we train DreamBooth on our text-to-video models. The quantitative and qualitative results are demonstrated in Tab. 1 and Fig. 6, where we can see that Dreambooth could hardly solve the identity mixing problem. Additionally, the tuning-based methods always require users to manually collect few-shot training samples for additional parameter-tuning, which not only can be time-consuming, but collecting same concept in different scenes as training samples is cumbersome, the situation is more intense for multi-concept video customization task. Therefore, our ConceptMaster is obviously more practical for real-world applications.

### v2

@@ Abstract
Text-to-video generation has made remarkable advancements through diffusion models. However, Multi-Concept Video Customization (MCVC) remains a significant challenge. We identify two key challenges for this task: 1) the identity decoupling issue, where directly adopting existing customization methods inevitably mix identity attributes when handling multiple concepts simultaneously, and 2) the scarcity of high-quality video-entity pairs, which is crucial for training a model that can well represent and decouple various customized concepts in video generation. To address these challenges, we introduce ConceptMaster, a novel framework that effectively addresses the identity decoupling issues while maintaining concept fidelity in video customization. Specifically, we propose to learn decoupled multi-concept embeddings and inject them into diffusion models in a standalone manner, which effectively guarantees the quality of customized videos with multiple identities, even for highly similar visual concepts. To overcome the scarcity of high-quality MCVC data, we establish a data construction pipeline, which enables collection of high-quality multi-concept video-entity data pairs across diverse scenarios. A multi-concept video evaluation set is further devised to comprehensively validate our method from three dimensions, including concept fidelity, identity decoupling ability, and video generation quality, across six different concept composition scenarios. Extensive experiments demonstrate that ConceptMaster significantly outperforms previous methods for video customization tasks, showing great potential to generate personalized and semantically accurate content for video diffusion models.
Figure 1: We propose ConceptMaster, a Multi-Concept Video Customization (MCVC) method that can create high-quality concept-consistent videos based on given multiple reference images without test-time tuning. Representatively, we demonstrate ConceptMaster’s video customization capacity on six scenarios, including 1) multiple persons, 2) persons with livings, 3) persons with stuffs, 4) multiple livings, 5) livings with stuffs and 6) persons with both livings and stuffs.
footnotetext: Work done during an internship at KwaiVGI, Kuaishou Technology Corresponding author
@@ 1 Introduction
Diffusion-based text-to-video generation models, trained on extensive text-video data pairs, have demonstrated remarkable success in generating high-quality videos from textual inputs [15, 6, 58, 3, 64, 56, 52, 42, 27, 19, 11, 2]. These advancements have sparked increasing interest in personalizing video generation through user-defined concepts. Recently, some methods have been proposed to produce customized videos using additional image guidance, with proven effectiveness for the customization on object [26], human [18], style [38], etc.
Existing approaches for concept customization primarily fall into two methodological categories: tuning-based solutions and pre-training-based methods. The tuning-based solutions [50, 13, 31, 16, 9] typically first optimize model parameters (e.g., variants of LoRAs [23] or fully training latent diffusion models [50]) each time for customized concepts and then incorporate them for inference. However, these methods are computationally time-consuming and often require the manual collection of multiple reference samples, rendering them impractical in most time-sensitive and user-friendly scenarios. Conversely, pre-training-based methods [62, 14, 34, 8, 67, 36, 60, 61, 26, 18] aim to integrate visual embeddings into diffusion models at training time in a data-driven manner, enabling personalization without additional test-time tuning. Despite their progress, how to adopt these approaches to simultaneously process multiple concepts in videos to keep both concept fidelity and factorization in a feed-forward approach remains challenging.
@@
To overcome these challenges, we propose ConceptMaster (see Fig. 1), an effective MCVC method that can effectively maintain the fidelity of multiple concepts and address the identity decoupling problem, even for highly similar concepts. Unlike previous approaches to integrate visual embeddings with textual counterparts [34, 36, 26], or directly aggregate visual embeddings as a whole into diffusion models [67, 61, 18], our key insight is to learn the decoupled multi-concept embeddings and inject into diffusion transformer models in a standalone manner. Specifically, the process includes: 1) Extracting comprehensive visual embeddings from given reference images, where we initially extract dense visual tokens by the CLIP image encoder [45], and integrate a learnable query transformer (Q-Former) network [35] to better represent comprehensive visual embeddings and align with the diffusion model space. 2) Incorporating visual representation with corresponding text description of every concept, where we propose the Decouple Attention Module (DAM) to conduct intra-pair attention to separately bind the extracted visual embedding with corresponding textual embedding for each concept, the process could effectively capture semantic differences across multiple concepts while maintain concept-specific uniqueness. 3) Introducing a novel multi-concept embeddings injection strategy, where we firstly composite the multi-concept embeddings and then inject them using an individual Multi-Concept Injector (MC-Injector), which is a standalone cross-attention layer, into the diffusion transformer models without affecting the original textual cross-attention. This strategy separates the functionality of the original textual cross-attention from the learning process of the newly injected composite multi-concept embeddings, effectively enhancing the representation of multiple identities. The designed ConceptMaster could efficiently create high-fidelity customized videos during inference without additional parameter tuning, which significantly provides the potential for the practicality of real-world applications.
Furthermore, to address the scarcity of suitable MCVC data, we carefully establish a data collection pipeline, which could collect high-quality MCVC data that precisely extract entity images and corresponding text descriptions of diverse concepts in videos. By utilizing this pipeline, we collect over 1.3 million video-entity pairs spanning diverse conceptual domains, including humans, livings, and various object categories. To further facilitate the evaluation, we introduce a multi-concept evaluation set to comprehensively validate this task from 1) concept fidelity, 2) effectiveness of identity decoupling, and 3) video generation quality across six distinct multi-concept composition scenes. Overall, our key contributions can be summarized:
@@
  • We present a novel strategy of learning decoupled multi-concept embeddings and injecting them into diffusion models without influencing the original attention operations, which effectively guarantees the fidelity of various concepts in customized videos.
  • We introduce a dedicated data construction pipeline that enables the collection of high-quality multi-concept video-entity pairs across diverse concepts, which effectively address the scarcity of high-quality MCVC data.
  • We collect a multi-concept evaluation set that could comprehensively validate video customization performance from six distinct concept composition scenarios and various dimensions including concept fidelity, identity decoupling and video quality. Extensive experiments demonstrate the superiority of ConceptMaster in video customization.
@@ 4 ConceptMaster
Figure 3: Overview of ConceptMaster framework. Given a caption along with a set of concept images and their semantic labels, we firstly extract comprehensive visual concept representations with the CLIP image encoder and a learnable Q-Former, then bind the visual representations with corresponding text embeddings of each concept through the Decouple Attention Module (DAM). Finally, the multi-concept visual-text embeddings are injected into the diffusion transformer models with the Multi-Concept Injector (MC-Injector).
@@ 4.1 Multi-Concept Video Customization
Problem: Given a caption $T$ describing a video, along with a set of concept images $\{X_{i}|i=1\ldots N\}$ and their corresponding labels $\{Y_{i}|i=1\ldots N\}$ (e.g., a man and a woman with their respective images), where $N$ represents the number of distinct concepts, the task of Multi-Concept Video Customization (MCVC) aims to generate high-quality videos that incorporate all image-defined visual concepts while aligning them with the given descriptive caption $T$. Each concept should maintain its identity as the provided images while precisely expressing its semantic behavior as described in the caption. We define the paire images and label for each concept as the intra-pair customized concept for convenience of expression.
@@
4.2 Decoupling and Injecting Concept Embeddings
@@
Decoupling Intra-Pair Embeddings. After obtaining the appropriate visual representation, we integrate the corresponding text labels to create visual-text aligned concept representation. While previous works [65, 36] directly combine the visual representation with the corresponding word from the caption embedding $c_{text}=\mathcal{E}_{\text{text}}(T)$, we hope to fully leverage the textual label information associated with related image to enhance the representation specific to each concept. Therefore, unlike these approaches, we employ T5-encoder $\mathcal{E}_{T5}$ to encode each concept label individually to obtain the text representation $\{y_{i}|y_{i}=\mathcal{E}_{T5}(Y_{i}),i=1\ldots N\}$. Subsequently, we introduce the Decouple Attention Module (DAM) to fuse each pair of the visual and text label embedding $\{(x_{i},y_{i})|i=1\ldots N\}$. The DAM operation can be formulated as:
{$&Q_{i}=W_{Q}\cdot x_{i};K_{i}=W_{K}\cdot y_{i};V_{i}=W_{V}\cdot y_{i},\\ &Attention(Q_{i},K_{i},V_{i})=\text{Softmax}(\frac{Q_{i}K_{i}^{T}}{\sqrt{d}})\cdot V_{i},\\ &b_{i}=Q_{i}+Attention(Q_{i},K_{i},V_{i})\\ &c_{i}=b_{i}+\text{FFN}(b_{i})\end{aligned}\right.$ (2)
where $W_{Q}$, $W_{K}$, and $W_{V}$ are projection matrices, $d$ is the embedding dimension, and FFN is a two-layer multi-layer perceptron (MLP) with GLUE [57] as the middle activation function. The residual connection [17] is existed in both attention and MLP layers. With the designed DAM, every visual representation could integrate its corresponding textual label to serve as the visual-text aligned representation for the diffusion transformer models.
@@ 5 Experiments
Methods Concept Fidelity and Decoupling Ability Video Quality
@@
CustomDiffusion 26.096 21.101 0.737 0.467 16.321 0.962 15.283 0.560 0.672
SSR-Encoder 23.766 21.153 0.739 0.466 16.651 0.946 21.804 0.535 0.657
IP-Adapter 26.097 21.980 0.749 0.492 16.425 0.969 17.800 0.496 0.693
MS-Diffusion 28.031 21.901 0.736 0.487 16.500 0.978 15.849 0.571 0.624
ConceptMaster (Ours) 28.246 22.165 0.781 0.584 16.169 0.967 26.115 0.572 0.657
Table 1: Quantitative comparison with different methods on the introduced multi-concept evaluation set, where bold represents the best result and underline represents the second best result.
@@ 5.1 Experimental Setup
Evaluation Metrics. In order to comprehensively evaluate MCVC methods, we consider three different dimensions: 1) Concept fidelity, where we exploit the commonly used CLIP-cap [45] to globally evaluate if the generated videos match the semantics of given video captions. 2) Decoupling ability, where we utilize LISA [33] to segment the mask area of each concept in generated videos, and then compute CLIP-I and DINO-I [4] scores between the original concept images and the mask areas of each concept in generated videos. Additionally, we compute the semantic matching scores between each concept label and its corresponding mask area. We refer CLIP-tag and CLIP-tagdis as the similarity and dissimilarity between label and its corresponding area. These four metrics are used to validate the model capacity for decoupling various concepts. 3) Video generation quality, where we adopt the motion smoothness, dynamic degree, aesthetic quality and imaging quality collected as suggested by VBench [25] to evaluate the generation quality.
Methods Concept Fidelity and Decoupling Ability Video Quality
@@
Merge Textual and Visual Embeddings 27.221 21.885 0.760 0.548 16.186 0.969 17.025 0.538 0.675
IP-Adapter-like 28.195 22.051 0.736 0.473 16.330 0.969 21.957 0.530 0.652
ConceptMaster (Ours) 28.246 22.165 0.781 0.584 16.169 0.967 26.115 0.572 0.657
Table 2: Quantitative comparison of different multi-concept embeddings injection manner, bold represents the best result and underline represents the second best result. All the methods we compared have been trained on the same data as that used by ConceptMaster.
Methods Concept Fidelity and Decoupling Ability Video Quality
@@
Without Q-Former 29.312 21.876 0.726 0.416 16.379 0.963 24.984 0.518 0.666
Without DAM 28.916 21.981 0.730 0.439 16.365 0.968 25.252 0.531 0.649
Concate-MLP 27.746 22.063 0.775 0.577 16.251 0.966 23.770 0.551 0.648
Self-Attn 28.146 22.045 0.778 0.576 16.211 0.968 22.864 0.550 0.655
DAM (Ours) 28.246 22.165 0.781 0.584 16.169 0.967 26.115 0.572 0.657
Table 3: Quantitative comparison of the design choice of Q-Former and DAM modules, bold represents the best result and underline represents the second best result. All the methods we compared have been trained on the same data as that used by ConceptMaster.
5.2 Multi-Concept Evaluation Set
In order to comprehensively evaluate the performance of MCVC methods, we establish a multi-concept evaluation set, including diverse concept composition scenarios of 1) multiple persons, 2) persons with livings, 3) persons with stuffs, 4) multiple livings, 5) livings with stuffs and 6) persons with both livings and stuffs. The sample number for each scenario is 40, 40, 40, 30, 30, and 30 respectively (Total 210). It should be noticed that we manually collect reference images and provide suitable captions for these scenarios, since we hope to eliminate information leakage for evaluation when extracting concepts from videos via the same MCVC data collection pipeline mentioned in Section 4.3 (e.g., the caption is A man and a woman are talking to each other, and the extracted man and woman are doing the same thing). All the quantitative experiments are evaluated by this introduced multi-concept evaluation set. More details are provided in supplementary material.
@@ 5.3 Comparing with other methods
We compare several open-sourced multi-concept image customization methods [31, 69, 67, 61], combining with the image-to-video (I2V) generation model I2VGen-XL [68], as a naive solution for the MCVC task with our ConceptMaster. According to the qualitative results in Fig. 5, we can see that our ConceptMaster has clear advantages on customizing multiple concepts in videos. The naive solution will be subject to the decoupling and representation ability of both two models, the instruction-following capability of I2V models could further influence the quality of generated videos. In contrast, the end-to-end video customization models could achieve better results. The quantitative results in Tab. 1 also demonstrates that our method could not only maintain the representation of multiple concepts, but also generate high-quality text-aligned videos.